# Probing new hadronic forces with heavy exotic atoms

### v1

@@ V Application to benchmark models
with
$g_{p(n)}^{\phi}=$ $\frac{m_{p(n)}}{f_{\phi}}\left(f_{u}^{p(n)}+f_{d}^{p(n)}+f_{s}^{p(n)}\right)\,,$ (13)
$g_{\pi(K)}^{\phi}=$ $-\frac{m_{\pi(K)}}{2f_{\phi}}\,,$ (14)
where $f^{p(n)}_{u}=0.018(0.016)$, $f^{p(n)}_{d}=0.034(0.038)$, and $f^{p(n)}_{s}=0.044(0.044)$ [108]. The interaction with nuclei is given by $g_{N}^{X}\approx(A-Z)g_{n}^{X}+Zg_{p}^{X}$.
The bounds on the $uds$-model are plotted in the left panel of Fig. 3 in the $m_{\phi}-f^{-1}_{\phi}\,(g_{p}^{\phi})$ plane, where we projected the bounds of Fig. 2. The approximated NLO bounds from rare kaon decay are $1/f_{\phi}\lesssim 0.8/\textrm{GeV}$ [72] and are not shown. We note that the probed range of parameter space is mildly fine-tuned, since for a TeV cutoff, the 1-loop correction to $m_{\pi}$ is $\delta m_{\phi}^{2}\sim\textrm{TeV}^{2}(m_{s}/f_{\phi})^{2}/16\pi^{2}$. To conclude, we learn that heavy exotic atoms already set the strongest bounds on the $uds$ model for $0.1\lesssim m_{\phi}\lesssim 10\,$ with the potential to improve by at least an order of magnitude.
The $B$-boson model introduces a new massive spin-1 gauge boson, $B$, with universal vector-like couplings to the quarks. The SM charged leptons are coupled via 1-loop kinetic mixing and we assume $\mathrm{BR}(B\to{\rm inv.})\approx 1$. This model was analyzed in several places, e.g. [110, 111]. The effective interactions with the SM fermions are given by
@@
where $q\,(\ell)$ is the SM quark (charged leptons). The interaction terms of Eq. (15) can be mapped to the hadron interactions
$\mathcal{L}_{B}^{\rm int,had}=g_{B}B_{\mu}\left(\bar{p}\gamma^{\mu}p+\bar{n}\gamma^{\mu}n\right),$ (16)
with no coupling to mesons. The baryon number is an anomalous symmetry, and thus requires additional degrees of freedom. Such considerations lead to strong bounds from rare bottom and kaon decays, see [112, 113, 114]. However, a similar model that is anomaly-free can be built by mixing the SM fermions with heavy vector-like fermions as in e.g. [115, 116].
We plot our new $\bar{p}\,$e and $\bar{p}\,$b bounds and the exotic atom projections for the $B$ model in the right panel of Fig. 3 (there are no bounds from mesonic atoms). We compare our result to the existing bounds from $\bar{p}^{4}$e, NA64 [117] and to rescaling of the $B-L$ bound from [17], denoted as CODATA. The last two bounds are based on a loop-induced $B$-electron coupling that depends on the model’s ultraviolet details and can be adjusted accordingly. We learn that $\bar{p}$e and $\bar{p}$b are already the robust leading probes of this model and that future measurements have the potential to increase sensitivity by an order of magnitude.
VI Conclusions
In this work, we explored the potential of precision spectroscopy of heavy exotic atoms in which electrons are replaced by $\bar{p}$, $\pi^{-}$ or $K^{-}$, to probe new force carriers with hadronic couplings. We have considered transitions that are clean from short-range hadronic effects, thus, the SM contribution can be predicted to high accuracy. However, nuclear polarizability can still have sizable effects on these transition lines. Therefore, we propose to use two lines and simultaneously solve for new-physics and polarizability effects. Based on the available data in $\bar{p}\,$e, $\bar{p}\,$b and $\pi^{-}\,$, we extract new world-leading bounds on two benchmark models and derive projections for future measurement for several representative systems. For mediator masses of up to 101010\,10, future experiments have the potential to improve the sensitivity by up to two orders of magnitude.
@@ Acknowledgements.
We thank N. Paul, K. Pachucki, M. Gorshteyn, and C. Delaunay for illuminating discussions and comments on the manuscript. Special thanks goes to P. Indelicato for valuable input on the existing experimental data and for testing prior calculations for consistency. The work of H.L. is supported by the U.S. Department of Energy under Grant Contract DE-SC0012704. The work of B.O. is supported by the Helen-Diller quantum center and ISF (grant no. 2071390). The work of O.S. and Y.S. is supported by grants from the NSF-BSF (grant No. 2021800), the ISF (grant No. 597/24).

### v2

@@ Abstract
with $g^{\phi}_{\pi(K)}=-\frac{m_{\pi(K)}}{2f_{\phi}}$, $g^{\phi}_{p(n)}=\frac{m_{p(n)}}{f_{\phi}}\sum_{q}f^{p(n)}_{q}$ and $f^{p(n)}_{u}=0.018(0.016)$, $f^{p(n)}_{d}=0.034(0.038)$, and $f^{p(n)}_{s}=0.044(0.044)$ [109]. The interaction with nuclei is given by $g_{N}^{X}\approx(A-Z)g_{n}^{X}+Zg_{p}^{X}$.
The bounds on the udsuds-model are plotted in the left panel of Fig. 3 in the $m_{\phi}-f^{-1}_{\phi}\,(g_{p}^{\phi})$ plane, where we projected the bounds of Fig. 2. The approximated NLO bounds from rare kaon decay are $1/f_{\phi}\lesssim 0.8/\textrm{GeV}$ [72] and are not shown. We note that the probed range of parameter space is mildly fine-tuned, since for a TeV cutoff, the 1-loop correction to $m_{\pi}$ is $\delta m_{\phi}^{2}\sim\textrm{TeV}^{2}(m_{s}/f_{\phi})^{2}/16\pi^{2}$. We learn that heavy exotic atoms already set the strongest bounds on the udsuds model for $0.1\lesssim m_{\phi}\lesssim 10\,\,\mathrm{MeV}$ with the potential to improve by at least an order of magnitude.
The $B$-boson model introduces a new massive spin-1 gauge boson, $B$, with universal vector-like couplings to the quarks. The SM charged leptons are coupled via 1-loop kinetic mixing and we assume $\mathrm{BR}(B\to{\rm inv.})\approx 1$. This model was analyzed in several places, e.g. [111, 112]. The interactions with the SM fermions are given by
@@
where $q\,(\ell)$ is the SM quarks (charged leptons). The $\mathcal{L}_{B}^{\rm int}$ can be mapped to the hadron interactions
$\mathcal{L}_{B}^{\rm int,had}=g_{B}B_{\mu}\left(\bar{p}\gamma^{\mu}p+\bar{n}\gamma^{\mu}n\right)\,.$ (14)
The baryon number is an anomalous symmetry, and thus requires additional degrees of freedom. Such considerations lead to strong bounds from rare bottom and kaon decays [113, 114, 115]. However, a similar model that is anomaly-free can be built by mixing the SM fermions with heavy vector-like fermions as in e.g. [116, 117].
We plot our new $\bar{p}\,$Ne and $\bar{p}\,$Pb bounds and the projections for the $B$ model in the right panel of Fig. 3. We compare our result to the existing bounds from $\bar{p}^{4}$He, NA64 [118] and to rescaling of the BLB-L bound from [17], denoted as CODATA. The last two are based on a loop-induced $B$-electron coupling that depends on the model’s UV details and can be adjusted accordingly. We learn that $\bar{p}$Ne and $\bar{p}$Pb are already the robust leading probes of this model and that future measurements have the potential to increase sensitivity by an order of magnitude.
Conclusions. We explored the potential of precision spectroscopy of heavy exotic atoms in which electrons are replaced by $\bar{p}$, $\pi^{-}$ or $K^{-}$, to probe new force carriers with hadronic couplings. We have considered transitions that are clean from short-range hadronic effects, thus, the SM contribution can be predicted to high accuracy. However, nuclear polarizability can still have sizable effects on these transition lines. Therefore, we propose to use two lines and simultaneously solve for new-physics and polarizability effects. Based on the available data in $\bar{p}\,$Ne, $\bar{p}\,$Pb and $\pi^{-}\,$N, we extract new world-leading bounds on two benchmark models and derive projections for future measurement for several representative systems. For mediator masses of up to $10\,\,\mathrm{MeV}$, future experiments have the potential to improve the sensitivity by up to two orders of magnitude. In addition, while this work considered only spin-independent interactions, spin-dependent interactions can be probed in a future work.
@@ Acknowledgements.
We thank N. Paul, K. Pachucki, M. Gorshteyn, and C. Delaunay for illuminating discussions and comments on the manuscript. Special thanks goes to P. Indelicato for valuable input on the existing experimental data and for testing prior calculations for consistency. The work of H.L. is supported by the U.S. Department of Energy under Grant Contract DE-SC0012704. The work of B.O. is supported by the Helen-Diller quantum center and ISF (grant no. 2071390). The work of O.S. and Y.S. is supported by grants from the NSF-BSF (grant No. 2021800), the ISF (grant No. 597/24). The work of O.S. is also supported by the Helen-Diller quantum center.